OpenStudy (anonymous):

solve: -7 is less than or equal to 3x-6 is less then or equal to 5 solution {xl

OpenStudy (a_clan):

\[-7\le3x-6\le5\] \[-1\le3x \le 11\] \[-1/3 \le x \le 11/3\]